OpenStudy (anonymous):

Find the equation of the line passing through the points (-2,-3) that is parallel to the line 5x-2y=10.

OpenStudy (anonymous):

y=(5/2)x+3

OpenStudy (anonymous):

Let's start with the piece of info that tells us that the line is parallel with 5x-2y = 10 cause that will give us the slope. 5x-2y = 10 -2y = -5x + 10 y = 5/2x + 2 from this we can see that our slope, m = 5/2. Now let's use the slope that we found and find the equation of the line that passes a given point (-2, -3) We know: y = mx +b -3 = (5/2) (-2) + b -3 = -5 + b b = 2 Therefore the equation of the line is y= 5/2x + 2

OpenStudy (anonymous):

where u get the 2 from?

OpenStudy (anonymous):

whenu divided the -2

OpenStudy (anonymous):

so the line is y=5/2x-5 find a line that has a slope of 5/2 and passes through (-2,-3) so the line is y=5/2x+2

OpenStudy (anonymous):

The second part?

OpenStudy (anonymous):

We were just solving for b ( the point where the line passes through the y -axis)